[6x + 5]/3 = 3x/7 – 2
Answer by stanbon(75887)

[6x + 5]/3 = 3x/7 – 2
Multiply thru by 21 to get:
7(6x+5) = 9x -42
42x+35 = 9x-42
33x = -77
3x = -7
x = -7/3
===============
Cheers,
Stan H.